# VAMO: Efficient Large-Scale Nonconvex Optimization via Adaptive Zeroth Order Variance Reduction

### v1

@@ Abstract
Optimizing large-scale nonconvex problems, common in machine learning, demands balancing rapid convergence with computational efficiency. First-order (FO) stochastic methods like SVRG provide fast convergence and good generalization but incur high costs due to full-batch gradients in large models. Conversely, zeroth-order (ZO) algorithms reduce this burden using estimated gradients, yet their slow convergence in high-dimensional settings limits practicality. We introduce VAMO (VAriance-reduced Mixed-gradient Optimizer), a stochastic variance-reduced method combining FO mini-batch gradients with lightweight ZO finite-difference probes under an SVRG-style framework. VAMO’s hybrid design uses a two-point ZO estimator to achieve a dimension-agnostic convergence rate of $\mathcal{O}(1/T+1/b)$, where $T$ is the number of iterations and $b$ is the batch-size, surpassing the dimension-dependent slowdown of purely ZO methods and significantly improving over SGD’s $\mathcal{O}(1/\sqrt{T})$ rate. Additionally, we propose a multi-point ZO variant that mitigates the $O(1/b)$ error by adjusting number of estimation points to balance convergence and cost, making it ideal for a whole range of computationally constrained scenarios. Experiments including traditional neural network training and LLM finetuning show VAMO outperforms established FO and ZO methods, offering a faster, more flexible option for improved efficiency.
@@ 1 Introduction
First-order (FO) optimization methods, particularly Stochastic Gradient Descent (SGD), have been applied in training a wide range of machine learning models. For large-scale problems, variance-reduced (VR) techniques, such as the Stochastic Variance Reduced Gradient (SVRG) algorithm allen2016improved ; reddi2016stochastic ; johnson2013accelerating , offer significant improvements, achieving faster convergence rates, $O(1/T)$, compared to the rate of SGD $O(1/\sqrt{T})$ reddi2016stochastic . However, in recent years, extremely large models—such as Large Language Models (LLMs) with billions of parameters—have become increasingly prevalent in machine learning. When training these models, traditional variance reduction methods like SVRG face a major challenge: they require periodically computing the full gradient over the entire dataset, which is often impractical for such large-scale models reddi2016stochastic . For LLMs, this step results in prohibitive computational and memory overhead, severely hindering efficient training of large models.
@@
Zeroth-order (ZO) optimization methods present an appealing alternative in this context, as they completely bypass the need for explicit gradient calculations, estimating gradients using only function value queries nesterov2017random ; liu2018zeroth . This gradient-free characteristic drastically reduces per-iteration computational cost and memory footprint, making ZO methods attractive for resource-constrained training of LLMs malladi2023fine ; gautam2024variance . Despite these advantages, ZO methods typically exhibit slower theoretical convergence rates than FO methods and, critically, often suffer from a strong dependence on the model dimension $d$ nesterov2017random ; duchi2015optimal . Given the vast dimensionality of modern LLMs, this dependence can render pure ZO approaches impractically slow. This creates a clear dilemma for large model training: FO methods offer desirable convergence, but suffer from high gradient costs, while ZO methods are cheaper per step but often too slow and scale poorly with model size. This naturally raises the question: can we devise a hybrid strategy that combines the strengths of both FO and ZO methods, thereby overcoming their individual limitations for efficient training of large models?
In this work, we propose VAMO (VAriance-reduced Mixed-gradient Optimizer), a new adaptive hybrid algorithm specifically designed to navigate this dilemma in large-scale non-convex optimization. Our approach integrates FO and ZO techniques within the SVRG framework, aiming to maintain SVRG’s fast convergence while substantially mitigating its computational burden. A major breakthrough here is the replacement of the prohibitively expensive full FO gradient $\nabla f(\hat{\mathbf{x}})$ at SVRG checkpoints with an efficient ZO gradient estimate ^$\hat{\nabla}f(\hat{\mathbf{x}})$, which significantly reduces computation. This leads to a convergence rate of $O(1/T+1/b)$, significantly outperforming both ZO methods and FO-SGD, matching the rate of FO-SVRG only with an additional complexity of $O(1/b)$, which we could further decrease by increasing computational budget. The adaptability of VAMO is then enhanced through several key innovations proposed in this work. First, we introduce a novel mixing coefficient into the update rule, which enables fine-grained control over the balance between the FO stochastic gradient and the ZO variance correction term. This mechanism allows practitioners to optimize performance based on the specific characteristics of the problem and the available computational budget. Secondly, we propose a configurable ZO gradient estimation strategy for the checkpoint gradient ^$\hat{\nabla}f(\hat{\mathbf{x}})$, which offers a choice between the standard two-point estimator and a more robust multi-point estimator. This flexibility introduces an additional degree of control, enabling users to balance the trade-off between estimation accuracy and the number of function evaluations. Together, these innovations make VAMO highly adaptable to a wide range of optimization scenarios. Crucially, despite this hybrid and adaptive nature, our gradient estimator is designed to maintain the unbiased property inherent to FO-SVRG, distinguishing our method from many biased ZO approaches and facilitating a rigorous convergence analysis for stronger theoretical guarantees.
@@ 2 Related Work
First-order optimization. While Stochastic Gradient Descent (SGD) robbins1951stochastic remains a foundational algorithm in machine learning, its convergence can be slow in large-scale settings due to high gradient variance johnson2013accelerating . Addressing this, variance-reduced (VR) methods gower2020variance , notably SVRG allen2016improved ; reddi2016stochastic ; johnson2013accelerating and SAGA defazio2014saga , represent a significant theoretical advancement. These algorithms reduce variance by leveraging gradients from past iterates or periodically computing full-batch gradients and achieve faster convergence rates (e.g., linear convergence under certain assumptions) compared to SGD reddi2016stochastic ; johnson2013accelerating . Despite these theoretical benefits, a primary practical limitation is the substantial computational and memory cost associated with full-batch gradients. This overhead can become prohibitive as model sizes and datasets scale. Another common extension of SGD is the use of adaptive step-sizes, as in ADAM kingma2014adam and Adagrad duchi2011adaptive . However, the convergence properties of these adaptive methods remain debated and can be highly sensitive to hyper-parameter choices reddi2019convergence ; defossez2020simple ; zhang2022adam .To provide a clearer and more interpretable comparison, we focus on standard baselines such as SGD and SVRG, which better isolate the effects of our proposed modifications. Notably, our algorithm can be readily combined with adaptive step-size techniques if needed.
@@
Zeroth-order optimization. Zeroth-order (ZO) optimization methods approximate gradients using function evaluations instead of explicit gradient computation, offering reduced computational and memory overhead zhang2024revisiting . This advantage makes them attractive for large-scale problems like large language model fine-tuning malladi2023fine ; gautam2024variance ; zhang2024revisiting ; tang2024private ; ling2024convergence , and their convergence properties are theoretically studied nesterov2017random ; liu2018zeroth ; duchi2015optimal ; jamieson2012query . However, ZO convergence rates often degrade with increasing problem dimension $d$, which makes them slow for high-dimensional models compared to FO methods liu2018zeroth ; liu2018stochastic ; wang2018stochastic . Even recent applications like MeZO malladi2023fine and MeZO-SVRG gautam2024variance are constrained by this dimensionality dependence. While ZO is crucial for black-box scenarios chen2017zoo ; tu2019autozoom , tasks like fine-tuning often have accessible gradients whose computation is merely expensive. This motivates our hybrid approach, which aims to combine ZO’s efficiency with FO’s faster convergence by strategically incorporating both types of gradient information, thereby avoiding the high computational cost associated with pure FO methods, while also mitigating the performance degradation that pure ZO methods often suffer in high-dimensional settings.
@@
Hybrid Zeroth-Order and First-Order Algorithms. Combining the strengths of FO and ZO optimization is a relatively recent and underexplored direction, with limited established theoretical analysis. The goal is to enjoy faster convergence while using less computational resource via ZO techniques zhang2024revisiting ; li2024addax . Early explorations include schemes like applying ZO to shallower model layers and FO to deeper ones zhang2024revisiting , or concurrently computing and combining FO-SGD and ZO-SGD updates at each step, as in Addax li2024addax . However, these initial hybrid approaches may have limitations; for instance, the theoretical convergence rate of Addax still exhibits dependence on the problem dimension $d$ li2024addax , hindering its effectiveness for large-scale models. Furthermore, many early hybrid strategies often lack explicit mechanisms to adaptively tune the balance between FO accuracy and ZO query efficiency in response to varying computational resources or specific problem demands. The scarcity of hybrid strategies that offer both theoretical convergence and controlled adaptability underscores the novelty of our work.
To contextualize our contributions, Table 1 summarizes the convergence rates and computational complexities of our proposed methods, referred to as VAMO and VAMO (multi-point) in the table alongside several FO and ZO algorithms. For ZO methods, ZO-SVRG is listed with a complexity of $O(nS+bT)$ function queries. Among FO methods, FO-SGD has the lowest computational cost $O(bdT)$ but also exhibits the slowest convergence ($O(1/\sqrt{T})$). FO-SVRG improves convergence to $O(1/T)$ but increases the cost to $O(ndS+bdT)$ due to full gradient computations. Our proposed VAMO maintains a complexity of $O(nS+bdT)$, similar to ZO-SVRG in terms of $nS$ but replacing the $ndS$ full gradient cost of FO-SVRG with a cheaper $nS$ ZO estimation cost for checkpoints, while achieving a fast $O(1/T+1/b)$ convergence rate. This makes its checkpoint cost significantly slower than FO-SVRG, especially when $d$ is large. The VAMO (multi-point) variant has a complexity of $O(qnS+bdT)$. Here, increasing $q$ (the number of ZO sampling directions) leads to higher complexity for checkpoint estimation but also improves the convergence rate to $O(1/T+\frac{1}{b}(1-\frac{q}{d})^{2})$, reducing the $O(1/b)$ error term and making its performance more comparable to FO-SVRG, particularly if $q\ll d$. This demonstrates that our proposed methods provide a flexible and often more efficient trade-off between computational cost and convergence performance compared to existing pure FO or ZO approaches. Our work further develops such an adaptive hybrid approach by specifically integrating ZO estimation within the SVRG structure, aiming to reduce the full-gradient cost while preserving strong convergence guarantees independent of dimensionality.
@@
Table 1: Summary of convergence rate and computational complexity of our proposals given $T$ total iterations. $n$ represents the total number of samples or component functions, $d$ is the problem dimension (number of parameters), $b$ denotes the mini-batch size, $S$ is the number of epochs or outer loops (for SVRG-type methods, $T\approx Sm$ where $m$ is the number of inner iterations per epoch), and $q$ signifies the number of query directions used for ZO estimation.
@@
ZO-SVRG Gradient Estimate $O\left(\frac{1}{d}\right)$ $O\left(\frac{d}{T}+\frac{1}{b}\right)$ $O(nS+bT)$
FO-SGD Explicit Gradient $O\left(\frac{1}{\sqrt{T}}\right)$ $O\left(\frac{1}{\sqrt{T}}\right)$ $O(bdT)$
FO-SVRG Explicit Gradient $O\left(1\right)$ $O\left(\frac{1}{T}\right)$ $O(ndS+bdT)$
VAMO Mixed Gradient $O\left(1\right)$ $O\left(\frac{1}{T}+\frac{1}{b}\right)$ $O(nS+bdT)$
VAMO(multi-point) Mixed Gradient $O\left(1\right)$ $O\left(\frac{1}{T}+\frac{1}{b}(1-\frac{q}{d})^{2}\right)$ $O(qnS+bdT)$
@@ 3 Preliminaries
where $\{f_{i}(\mathbf{x})\}_{i=1}^{n}$ are $n$ individual nonconvex cost functions. Note that Eq. (1) is the generic form of many machine learning problems such as training neural networks, since this is the natural form arising from empirical risk minimization (ERM). Next we introduce assumptions we will make throughout the paper and provide the background of ZO gradient estimate.
@@ 3.1 Assumptions
These assumptions are standard in the analysis of stochastic optimization algorithms for nonconvex problems reddi2016stochastic ; nesterov2017random ; ghadimi2013stochastic .
@@ 3.2 Convergence Notion
This work addresses the nonconvex optimization problem defined in Eq. (1), a setting prevalent in modern machine learning, particularly deep learning. In convex problems where local minima are global, …the convergence is typically measured by the expected suboptimality $\mathbb{E}\left[f(\mathbf{x}_{T})-f(\mathbf{x}^{*})\right]$. However, in nonconvex settings, identifying a global minimum is generally intractable due to the potential presence of multiple local minima and saddle points dauphin2014identifying ; balasubramanian2022zeroth .
Consequently, for such nonconvex problems, the convergence is evaluated by the first-order stationary condition in terms of the expected squared gradient norm $\mathbb{E}[\|\nabla f(\mathbf{x})\|_{2}^{2}]$. An algorithm is considered to converge if this metric approaches zero or falls below a specified tolerance $\epsilon$ liu2020primer ; mu2024variance . It serves as the primary metric for the theoretical convergence guarantees presented in this paper.
@@ 3.3 ZO Gradient Estimation
Consider an individual cost function $f_{i}:\mathbb{R}^{d}\to\mathbb{R}$ that satisfies the conditions in Assumption 1. The ZO approach estimates gradients using only function evaluations.
A commonly used two-point ZO gradient estimator for $f_{i}(\mathbf{x})$ is defined as nesterov2017random ; spall1992multivariate :
^$\hat{\nabla}f_{i}(\mathbf{x})=\frac{d}{\mu}\left[f_{i}(\mathbf{x}+\mu\mathbf{u}_{i})-f_{i}(\mathbf{x})\right]\mathbf{u}_{i},\quad\text{for }i\in[n],$ (2)
where $d$ is the dimensionality of the parameter vector $\mathbf{x}$, $\mu>0$ is a small smoothing parameter, and $\{\mathbf{u}_{i}\}_{i=1}^{n}$ are i.i.d. random direction vectors drawn uniformly from the unit Euclidean sphere in $\mathbb{R}^{d}$ (i.e., $\mathbf{u}_{i}\sim U(\mathbb{S}^{d-1})$) flaxman2004online ; shamir2017optimal ; gao2018information .
In general, for $\mu>0$, the ZO gradient estimator ^$\hat{\nabla}f_{i}(\mathbf{x})$ is a biased approximation of the true gradient $\nabla f_{i}(\mathbf{x})$. The bias tends to decrease as $\mu\to 0$. However, in practical implementations, choosing an excessively small $\mu$ can render the function difference $f_{i}(\mathbf{x}+\mu\mathbf{u}_{i})-f_{i}(\mathbf{x})$ highly sensitive to numerical errors or system noise or numerical precision issues, potentially failing to accurately represent the local change in the function lian2016comprehensive . A key property of the ZO estimator is that for $\mu>0$, it provides an unbiased estimate of the gradient of a smoothed version of $f_{i}$, often denoted $f_{i,\mu}(\mathbf{x})=\mathbb{E}_{\mathbf{v}}[f_{i}(\mathbf{x}+\mu\mathbf{v})]$ (where $\mathbf{v}$ is a random vector from a unit ball or sphere), i.e., $\mathbb{E}_{\mathbf{u}_{i}}[\hat{\nabla}f_{i}(\mathbf{x})]=\nabla f_{i,\mu}(\mathbf{x})$ liu2020primer .
To reduce the variance of the ZO gradient estimate, a multi-point version can be employed. Instead of using a single random direction $\mathbf{u}_{i}$, $q\geq 1$ i.i.d. random directions $\{\mathbf{u}_{i,j}\}_{j=1}^{q}$ are sampled for each $f_{i}$. Since estimating along each direction requires two function queries, the multi-point ZO gradient estimator involves a total of $2q$ function queries and is defined as duchi2015optimal ; liu2020primer :
^$\hat{\nabla}f_{i}(\mathbf{x})=\frac{d}{\mu q}\sum_{j=1}^{q}\left[f_{i}(\mathbf{x}+\mu\mathbf{u}_{i,j})-f_{i}(\mathbf{x})\right]\mathbf{u}_{i,j},\quad\text{for }i\in[n].$ (3)

### v2

# VAMO: Efficient Zeroth-Order Variance Reduction for SGD with Faster Convergence

@@ Abstract
Optimizing large-scale nonconvex problems, common in deep learning, demands balancing rapid convergence with computational efficiency. First-order (FO) optimizers, which serve as today’s baselines, provide fast convergence and good generalization but often incur high computation and memory costs due to the large size of modern models. Conversely, zeroth-order (ZO) algorithms reduce this burden using estimated gradients, yet their slow convergence in high-dimensional settings limits practicality. We introduce VAMO (VAriance-reduced Mixed-gradient Optimizer), a stochastic variance-reduced method that extends mini-batch SGD with full-batch ZO gradients under an SVRG-style framework. VAMO’s hybrid design utilizes a two-point ZO estimator to achieve a dimension-agnostic convergence rate of $\mathcal{O}(1/T+1/b)$, where $T$ is the number of iterations and $b$ is the batch-size, surpassing the dimension-dependent slowdown of purely ZO methods and significantly improving over SGD’s $\mathcal{O}(1/\sqrt{T})$ rate. Additionally, we propose a multi-point variant that mitigates the $O(1/b)$ error by adjusting the number of estimation points to balance convergence and cost. Importantly, VAMO achieves these gains with smaller dynamic memory requirements than many FO baselines, making it particularly attractive for edge deployment. Experiments including traditional neural network training and LLM finetuning confirm that VAMO not only outperforms established FO and ZO methods, but also does so with a light memory footprint.
@@ 1 Introduction
First-order (FO) optimization methods, particularly Stochastic Gradient Descent (SGD), have been applied in training a wide range of machine learning models. For large-scale problems, variance-reduced (VR) techniques, such as the Stochastic Variance Reduced Gradient (SVRG) algorithm (Johnson & Zhang, 2013; Allen-Zhu & Yuan, 2016; Reddi et al., 2016), offer significant improvements, achieving faster convergence rates, $O(1/T)$, compared to the rate of SGD $O(1/\sqrt{T})$ (Reddi et al., 2016). However, in recent years, extremely large models—such as Large Language Models (LLMs) with billions of parameters—have become increasingly prevalent in machine learning. When training these models, traditional variance reduction methods like SVRG face a major challenge: they require periodically computing the full gradient over the entire dataset, which requires high cost for such large-scale models (Reddi et al., 2016). For LLMs, this step results in prohibitive computational and memory overhead, severely hindering efficient training of large models.
@@
Zeroth-order (ZO) optimization methods present an appealing alternative in this context, as they completely bypass the need for explicit gradient calculations, estimating gradients using only function value queries (Nesterov & Spokoiny, 2017; Liu et al., 2018b). This gradient-free characteristic drastically reduces per-iteration computational cost and memory footprint, making ZO methods attractive for resource-constrained training of LLMs (Malladi et al., 2023; Gautam et al., 2024). Despite these advantages, ZO methods typically exhibit slower theoretical convergence rates than FO methods and, critically, often suffer from a strong dependence on the parameter dimension $d$ (Duchi et al., 2015; Nesterov & Spokoiny, 2017). Given the vast dimensionality of modern LLMs, this dependence can render pure ZO approaches impractically slow. This creates a clear dilemma for large model training: FO methods offer desirable convergence, but suffer from high gradient costs, while ZO methods are cheaper per step but often too slow and scale poorly with the parameter dimension. This naturally raises the question: can we devise a hybrid strategy that combines the strengths of both FO and ZO methods, thereby overcoming their limitations for efficient training of large models?
[…]
@@ 2 Related Work
While Stochastic Gradient Descent (SGD) (Robbins & Monro, 1951) remains a foundational algorithm in machine learning, its convergence can be slow in large-scale settings due to high gradient variance (Johnson & Zhang, 2013). Addressing this, variance-reduced (VR) methods (Gower et al., 2020), notably SVRG (Johnson & Zhang, 2013; Allen-Zhu & Yuan, 2016; Reddi et al., 2016) and SAGA (Defazio et al., 2014), represent a significant theoretical advancement. These algorithms reduce variance by leveraging gradients from past iterates or periodically computing full-batch gradients and achieve faster convergence rates (e.g., linear convergence under certain assumptions) compared to SGD (Johnson & Zhang, 2013; Reddi et al., 2016). Despite these theoretical benefits, a primary practical limitation is the substantial computational and memory cost associated with full-batch gradients. This overhead can become prohibitive as model sizes and datasets scale. Another common extension of SGD is the use of adaptive step-sizes, as in Adagrad (Duchi et al., 2011) and ADAM (Kingma & Ba, 2014). However, the convergence properties of these adaptive methods remain debated and can be highly sensitive to hyper-parameter choices (Reddi et al., 2019; Défossez et al., 2020; Zhang et al., 2022). To provide a clearer and more interpretable comparison, we focus on standard baselines such as SGD and SVRG on our theoretical analysis, which better isolate the effects of our proposed modifications.
Zeroth-order optimization.
@@
Zeroth-order (ZO) optimization approximates gradients using function evaluations instead of explicit gradient computation, offering reduced computational and memory overhead (Zhang et al., 2024). This advantage makes them attractive for large-scale problems like fine-tuning of large language models (Malladi et al., 2023; Gautam et al., 2024; Zhang et al., 2024; Tang et al., 2024; Ling et al., 2024), and their convergence properties are theoretically studied (Duchi et al., 2015; Jamieson et al., 2012; Nesterov & Spokoiny, 2017; Liu et al., 2018b). However, the convergence rates of ZO methods often degrade with increasing parameter dimension $d$, which makes them slow for large models compared to FO methods (Liu et al., 2018b; a; Wang et al., 2018). Even recent applications like MeZO (Malladi et al., 2023) and MeZO-SVRG (Gautam et al., 2024) are constrained by large parameter dimension. While ZO optimization is crucial for black-box scenarios (Chen et al., 2017; Tu et al., 2019), tasks like fine-tuning often have accessible gradients whose computation is merely expensive. This motivates our hybrid approach, which aims to combine ZO’s efficiency with FO’s faster convergence by strategically incorporating both types of gradient information, thereby avoiding the high computational cost associated with FO methods, while also mitigating the performance degradation that ZO methods often suffer in high-dimensional settings.
Hybrid Zeroth-Order and First-Order Algorithms.
@@
Combining the strengths of FO and ZO optimization is a relatively recent and underexplored direction, with limited established theoretical analysis. The goal is to enjoy faster convergence while using less computational resource via ZO techniques (Zhang et al., 2024; Li et al., 2024). Early explorations include schemes like applying ZO to shallower model layers and FO to deeper ones (Zhang et al., 2024), or concurrently computing and combining FO-SGD and ZO-SGD updates at each step, as in Addax (Li et al., 2024). However, these initial hybrid approaches may have limitations; for instance, the theoretical convergence rate of Addax still exhibits dependence on the problem dimension $d$ (Li et al., 2024), hindering its effectiveness for large-scale models. Furthermore, many early hybrid strategies often lack explicit mechanisms to adaptively tune the balance between FO accuracy and ZO query efficiency in response to varying computational resources or specific problem demands. The scarcity of hybrid strategies that offer both theoretical convergence and controlled adaptability underscores the novelty of our work. To contextualize our contributions, Table 1 summarizes the convergence rates and computational complexities of our proposed methods, referred to as VAMO and VAMO (multi-point) in the table alongside several FO and ZO algorithms. We provide a detailed explanation of Table 1 in Appendix B.
@@
Table 1: Summary of convergence rate and computational complexity of our proposals given $T$ total iterations. $n$ represents the total number of samples or component functions, $d$ is the parameter dimension, $b$ denotes the mini-batch size, $S$ is the number of epochs or outer loops (for SVRG-type methods, $T\approx Sm$ where $m$ is the number of inner iterations per epoch), and $q$ signifies the number of query directions used for ZO estimation.
@@
ZO-SVRG Gradient Estimate $O(1/d)$ $O(d/T+1/b)$ $O(nS+bT)$
FO-SGD Explicit Gradient $O(1/\sqrt{T})$ $O(1/\sqrt{T})$ $O(bdT)$
FO-SVRG Explicit Gradient $O(1)$ $O(1/T)$ $O(dnS+bdT)$
VAMO Mixed Gradient $O(1)$ $O(1/T+1/b)$ $O(nS+bdT)$
VAMO(multi-point) Mixed Gradient $O(1)$ $O(1/T+(1-q/d)^{2}/b)$ $O(qnS+bdT)$
@@ 3 Preliminaries
where $\{f_{i}(\mathbf{x})\}_{i=1}^{n}$ are $n$ individual nonconvex cost functions. Note that  equation 1 is the generic form of many machine learning problems such as training neural networks, since this is the natural form arising from empirical risk minimization (ERM). Next we introduce assumptions we will make throughout the paper and provide the background of ZO gradient estimate.
@@ 3.1 Assumptions
These assumptions are standard in the analysis of stochastic optimization algorithms for nonconvex problems (Ghadimi & Lan, 2013; Reddi et al., 2016; Nesterov & Spokoiny, 2017).
@@ 3.2 Convergence Notion
We study the nonconvex optimization problem in equation 1, which is common in modern machine learning. For convex problems, convergence is typically measured by the expected suboptimality $\mathbb{E}[f(\mathbf{x}_{T})-f(\mathbf{x}^{*})]$. In contrast, nonconvex problems often contain multiple local minima and saddle points (Dauphin et al., 2014; Balasubramanian & Ghadimi, 2022), making global optimality intractable. Consequently, convergence is evaluated by the first-order stationarity condition via the expected squared gradient norm $\mathbb{E}[\|\nabla f(\mathbf{x})\|_{2}^{2}]$. An algorithm is deemed convergent when this metric approaches zero or falls below a tolerance $\epsilon$ (Liu et al., 2020; Mu et al., 2024), which is the criterion used in our theoretical guarantees.
@@ 3.3 ZO Gradient Estimation
Consider an individual cost function $f_{i}:\mathbb{R}^{d}\to\mathbb{R}$ that satisfies the conditions in Assumption 1. The ZO approach estimates gradients using only function evaluations. A commonly used two-point ZO gradient estimator for $f_{i}(\mathbf{x})$ is defined as (Spall, 1992; Nesterov & Spokoiny, 2017):
$\hat{\nabla}f_{i}(\mathbf{x})=\frac{d}{\mu}\left[f_{i}(\mathbf{x}+\mu\mathbf{u}_{i})-f_{i}(\mathbf{x})\right]\mathbf{u}_{i},\quad\text{for }i\in[n],$ (2)
where $d$ is the dimensionality of the parameter vector $\mathbf{x}$, $\mu>0$ is a small smoothing parameter, and $\{\mathbf{u}_{i}\}_{i=1}^{n}$ are i.i.d. random direction vectors drawn uniformly from the unit Euclidean sphere in $\mathbb{R}^{d}$ (i.e., $\mathbf{u}_{i}\sim U(\mathbb{S}^{d-1})$) (Flaxman et al., 2004; Shamir, 2017; Gao et al., 2018).
In general, for $\mu>0$, the ZO gradient estimator $\hat{\nabla}f_{i}(\mathbf{x})$ is a biased approximation of the true gradient $\nabla f_{i}(\mathbf{x})$. The bias tends to decrease as $\mu\to 0$. However, in practical implementations, choosing an excessively small $\mu$ can render the function difference $f_{i}(\mathbf{x}+\mu\mathbf{u}_{i})-f_{i}(\mathbf{x})$ highly sensitive to numerical errors or system noise or numerical precision issues, potentially failing to accurately represent the local change in the function (Lian et al., 2016). A key property of the ZO estimator is that for $\mu>0$, it provides an unbiased estimate of the gradient of a smoothed version of $f_{i}$, often denoted $f_{i,\mu}(\mathbf{x})=\mathbb{E}_{\mathbf{v}}[f_{i}(\mathbf{x}+\mu\mathbf{v})]$ (where $\mathbf{v}$ is a random vector from a unit ball or sphere), i.e., $\mathbb{E}_{\mathbf{u}_{i}}[\hat{\nabla}f_{i}(\mathbf{x})]=\nabla f_{i,\mu}(\mathbf{x})$ (Liu et al., 2020).
To reduce the error of the ZO gradient estimate, a multi-point version can be employed. Instead of using a single random direction $\mathbf{u}_{i}$, $q\geq 1$ i.i.d. random directions $\{\mathbf{u}_{i,j}\}_{j=1}^{q}$ are sampled for each $f_{i}$. Since estimating along each direction requires two function queries, the multi-point ZO gradient estimator involves a total of $2q$ function queries and is defined as (Duchi et al., 2015; Liu et al., 2020):
$\hat{\nabla}f_{i}(\mathbf{x})=\frac{d}{\mu q}\sum_{j=1}^{q}\left[f_{i}(\mathbf{x}+\mu\mathbf{u}_{i,j})-f_{i}(\mathbf{x})\right]\mathbf{u}_{i,j},\quad\text{for }i\in[n].$ (3)